Astrid1:

Idk how to solve this -.- −4x − 2y = −12 4x + 8y = −24

babygirlarie:

what are you looking for ...x or y

Astrid1:

Im not sure, It says that we have to solve each equation by elimination.

xxDeppressionxx:

You are looking for both x and y

xxDeppressionxx:

so you would do this Solve for xx in -4x-2y=-12. x= 2/6−y ​ Substitute x=2/6−y ​ into 4x+8y=-24. 12+6y=-24 Solve for y in 12+6y=-24. y=−6 Therefore, x=6 y=−6 ​

babygirlarie:

add them together 4x-4x-8y-2y=-24-12 0x-10y=-36 -10y=-36 10y=36 divide 10 y=3.6 subsitute for 3.6 for y 4x-8(3.6)=-24 4x-28.8=-24 add 28.8 to both sdies 4x=4.8 divide 4 x=1.2 x=1.2 y=3.6

Astrid1:

@xxDeppressionxx how did you get X=2/6-y?

xxDeppressionxx:

I don't remember i'm sorry

xxDeppressionxx:

i can try to thiough

xxDeppressionxx:

though*

Astrid1:

~~Okie~~ o: Please try and remember o:

Aidsinyou:

Magic tbh

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aidsinyou Magic tbh \(\color{#0cbb34}{\text{End of Quote}}\) -.- Dats not gonna hewp me o:

Aidsinyou:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Aidsinyou Magic tbh \(\color{#0cbb34}{\text{End of Quote}}\) -.- Dats not gonna hewp me o: \(\color{#0cbb34}{\text{End of Quote}}\) If you learn sorcery it would help.

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aidsinyou \(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Aidsinyou Magic tbh \(\color{#0cbb34}{\text{End of Quote}}\) -.- Dats not gonna hewp me o: \(\color{#0cbb34}{\text{End of Quote}}\) If you learn sorcery it would help. \(\color{#0cbb34}{\text{End of Quote}}\) Id wanna learn Sorcery -.-

Aidsinyou:

This is a good calculator it shows how to do it as well

Aidsinyou:

uwu

Astrid1:

I know both of those calculators and it may show me how to use it but i want to actually actually know how to do it.

Aidsinyou:

Oh, well I forgot how to solve those, sorry Astrid

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aidsinyou Oh, well I forgot how to solve those, sorry Astrid \(\color{#0cbb34}{\text{End of Quote}}\) Its okay!

xxDeppressionxx:

Slope = -4.000/2.000 = -2.000 x-intercept = 6/2 = 3 y-intercept = 6/1 = 6.00000 for the first one

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Slope = -4.000/2.000 = -2.000 x-intercept = 6/2 = 3 y-intercept = 6/1 = 6.00000 for the first one \(\color{#0cbb34}{\text{End of Quote}}\) Ah! Ty!

xxDeppressionxx:

Slope = -1.000/2.000 = -0.500 x-intercept = -6/1 = -6.00000 y-intercept = -6/2 = -3 for the second one